1. Two charges Q1=-5 NC and Q2=-7 nc are separated by 8 cm. Calculate tHe electric potential at point.

2. A research van de Graaf generator has a 2.50 mm diameter sphere with a charge of 6.00 UC on it.

A. What is the potential near its surface?

B. At what distance from its center is the potential 1.50 MV?


1
Expert's answer
2022-03-02T14:39:37-0500

1.

"\varphi=\frac kr(q_1-q_2)=225~V."

2.

A.

"\varphi=\frac{kq}r=21.6~MV,"

B.

"r=\frac{kq}{\varphi}=36~mm."